# An Accelerated Algorithm for Stochastic Bilevel Optimization under Unbounded Smoothness

### v4

@@ 6 Conclusion
In this paper, we propose a new algorithm named AccBO for solving bilevel optimization problems where the upper-level is nonconvex and unbounded smooth and the lower-level problem is strongly convex. The algorithm achieved $\widetilde{O}(\epsilon^{-3})$ oracle complexity for finding an $\epsilon$-stationary point, which matches the rate of the state-of-the-art single-level relaxed smooth optimization [53] and is nearly optimal in terms of dependency on $\epsilon$ [2].
Limitations. One limitation of our work is that the convergence analysis for the Option I of our algorithm relies on the lower-level problem being a quadratic function: only under this case the algorithm becomes a single-loop procedure. However, it remains unclear how to design single-loop algorithms for more general lower-level strongly convex functions.
@@
We would like to thank the anonymous reviewers for their helpful comments. We would like to thank Tianbao Yang for improving our proof for the quadratic function with distributional drift in the earlier version of our paper. This work has been supported by the Presidential Scholarship, the ORIEI seed funding, and the IDIA P3 fellowship from George Mason University, the Cisco Faculty Research Award, and NSF award #2436217, #2425687. The Computations were run on ARGO, a research computing cluster provided by the Office of Research Computing at George Mason University (URL: https://orc.gmu.edu).
@@ Appendix A Technical Lemmas
In this section, we will introduce a few useful lemmas. The following technical lemma on recursive control is crucial for providing high probability guarantee of the lower-level variables $y_{t}$ and $\hat{y}_{t}$ in Algorithm 2 at anytime. We follow a similar argument as in [18, Proposition 29] with a slight generalization.
@@ Appendix B Auxiliary Lemmas for Bilevel Optimization
Lemma B.1 (Lipschitz property, [36, Lemma 8]).
@@
Lemma B.2 ($(L_{0},L_{1})$-smoothness, [36, Lemma 9]).
@@
Lemma B.3 (Descent inequality, [36, Lemma 10]).
@@
$\left.\quad\quad-\frac{2\mu\sqrt{2\alpha}}{1+\sqrt{\mu\alpha}}\alpha(1-L\alpha)\langle u_{t,2},\varepsilon_{t}\rangle\frac{1+\sqrt{\mu\alpha}}{2\sqrt{2\alpha}}\|z_{t}-w_{t}\|+\frac{L\alpha^{2}}{2}\|\varepsilon_{t}\|^{2}\right]+\frac{20\mu\Delta_{t}^{2}}{\sqrt{\mu\alpha}}.$
@@
Let us define vector $\omega_{t}\in\mathbb{R}^{2d}$ and matrices ${\mathbf{A}},{\mathbf{B}}\in\mathbb{R}^{2d\times 2d}$ as
$\omega_{t}=\begin{bmatrix}\nabla\phi_{t}(z_{t})\\ \varepsilon_{t}\end{bmatrix},\quad{\mathbf{A}}=\begin{bmatrix}1+\gamma&-\gamma\\ 1&0\end{bmatrix}\otimes{\mathbf{I}}_{d},\quad{\mathbf{B}}=\begin{bmatrix}-\alpha&-\alpha\\ 0&0\end{bmatrix}\otimes{\mathbf{I}}_{d}.$
By (9) we have $[(w_{t+1}-w_{t}^{*})^{\top},(w_{t}-w_{t}^{*})^{\top}]^{\top}={\mathbf{A}}\theta_{t}+{\mathbf{B}}\omega_{t}$. Since $\phi_{t}$ is $\mu$-strongly convex, then
@@
where matrix ${\mathbf{X}}_{1}\in\mathbb{R}^{4d\times 4d}$ is defined as
${\mathbf{X}}_{1}=\frac{1}{2}\begin{bmatrix}\mu\gamma^{2}&-\mu\gamma^{2}&-\gamma&0\\ -\mu\gamma^{2}&\mu\gamma^{2}&\gamma&0\\ -\gamma&\gamma&\alpha(2-L\alpha)&0\\ 0&0&0&0\end{bmatrix}\otimes{\mathbf{I}}_{d}.$
@@
where matrix ${\mathbf{X}}_{2}\in\mathbb{R}^{4d\times 4d}$ is defined as
${\mathbf{X}}_{2}=\frac{1}{2}\begin{bmatrix}\mu(1+\gamma)^{2}&-\mu\gamma(1+\gamma)&-(1+\gamma)&0\\ -\mu\gamma(1+\gamma)&\mu\gamma^{2}&\gamma&0\\ -(1+\gamma)&\gamma&\alpha(2-L\alpha)&0\\ 0&0&0&0\end{bmatrix}\otimes{\mathbf{I}}_{d}.$
@@
With definition of $\gamma,\rho^{2}$ and ${\mathbf{P}}$ given in the statement of the theorem, by direct calculation we have
$\begin{bmatrix}{\mathbf{A}}^{\top}{\mathbf{P}}{\mathbf{A}}-\rho^{2}{\mathbf{P}}&{\mathbf{A}}^{\top}{\mathbf{P}}{\mathbf{B}}\\ {\mathbf{B}}^{\top}{\mathbf{P}}{\mathbf{A}}&{\mathbf{B}}^{\top}{\mathbf{P}}{\mathbf{B}}\end{bmatrix}-(\rho^{2}{\mathbf{X}}_{1}+(1-\rho^{2}){\mathbf{X}}_{2})$
$\quad\quad\quad\quad\quad=\begin{bmatrix}-\frac{\mu(1-\sqrt{\mu\alpha})^{3}}{2\sqrt{\mu\alpha}(1+\sqrt{\mu\alpha})}&\frac{\mu(1-\sqrt{\mu\alpha})^{3}}{2\sqrt{\mu\alpha}(1+\sqrt{\mu\alpha})}&0&-\frac{1+\mu\alpha}{2(1+\sqrt{\mu\alpha})}\\ \frac{\mu(1-\sqrt{\mu\alpha})^{3}}{2\sqrt{\mu\alpha}(1+\sqrt{\mu\alpha})}&-\frac{\mu(1-\sqrt{\mu\alpha})^{3}}{2\sqrt{\mu\alpha}(1+\sqrt{\mu\alpha})}&0&\frac{1-\sqrt{\mu\alpha}}{2(1+\sqrt{\mu\alpha})}\\ 0&0&-\frac{\alpha(1-L\alpha)}{2}&\frac{\alpha}{2}\\ -\frac{1+\mu\alpha}{2(1+\sqrt{\mu\alpha})}&\frac{1-\sqrt{\mu\alpha}}{2(1+\sqrt{\mu\alpha})}&\frac{\alpha}{2}&\frac{\alpha}{2}\end{bmatrix}\otimes{\mathbf{I}}_{d}:={\mathbf{C}},$
where matrix ${\mathbf{C}}$ is negative semidefinite [12], i.e., ${\mathbf{C}}\preceq\mathbf{0}$. Then we use this fact to leverage
$\begin{bmatrix}w_{t+1}-w_{t}^{*}\\ w_{t}-w_{t}^{*}\end{bmatrix}^{\top}{\mathbf{P}}\begin{bmatrix}w_{t+1}-w_{t}^{*}\\ w_{t}-w_{t}^{*}\end{bmatrix}-\rho^{2}\theta_{t}^{\top}{\mathbf{P}}\theta_{t}=\begin{bmatrix}\theta_{t}\\ \omega_{t}\end{bmatrix}^{\top}(\rho^{2}{\mathbf{X}}_{1}+(1-\rho^{2}){\mathbf{X}}_{2}+{\mathbf{C}})\begin{bmatrix}\theta_{t}\\ \omega_{t}\end{bmatrix}$
$\quad\quad\leq\begin{bmatrix}\theta_{t}\\ \omega_{t}\end{bmatrix}^{\top}(\rho^{2}{\mathbf{X}}_{1}+(1-\rho^{2}){\mathbf{X}}_{2})\begin{bmatrix}\theta_{t}\\ \omega_{t}\end{bmatrix}$
$\quad\quad\leq-(\phi_{t}(w_{t+1})-\phi_{t}(w_{t}^{*}))+\rho^{2}(\phi_{t}(w_{t})-\phi_{t}(w_{t}^{*}))-\alpha(1-L\alpha)\langle\nabla\phi_{t}(z_{t}),\varepsilon_{t}\rangle+\frac{L\alpha^{2}}{2}\|\varepsilon_{t}\|^{2},$
where the last inequality follows by (19). Rearrange the above inequality and by definition of the potential function $V_{t}$, we obtain
@@
$\quad\quad\leq\rho^{2}V_{t}-\mu\alpha(1-L\alpha)\langle z_{t}-w_{t}^{*},\varepsilon_{t}\rangle+\frac{L\alpha^{2}}{2}\|\varepsilon_{t}\|^{2}$
$\quad\quad=\rho^{2}V_{t}-\mu\alpha(1-L\alpha)\langle w_{t}-w_{t}^{*},\varepsilon_{t}\rangle-\mu\alpha(1-L\alpha)\langle z_{t}-w_{t},\varepsilon_{t}\rangle+\frac{L\alpha^{2}}{2}\|\varepsilon_{t}\|^{2}$
@@
$\quad\quad\quad\quad\quad\quad-\frac{2\mu\sqrt{2\alpha}}{1+\sqrt{\mu\alpha}}\alpha(1-L\alpha)\langle u_{t,2},\varepsilon_{t}\rangle\frac{1+\sqrt{\mu\alpha}}{2\sqrt{2\alpha}}\|z_{t}-w_{t}\|+\frac{L\alpha^{2}}{2}\|\varepsilon_{t}\|^{2},$
@@
$\left.\quad\quad\quad\quad\quad\quad\quad-\frac{2\mu\sqrt{2\alpha}}{1+\sqrt{\mu\alpha}}\alpha(1-L\alpha)\langle u_{t,2},\varepsilon_{t}\rangle\frac{1+\sqrt{\mu\alpha}}{2\sqrt{2\alpha}}\|z_{t}-w_{t}\|+\frac{L\alpha^{2}}{2}\|\varepsilon_{t}\|^{2}\right]$
@@
$\left.\quad\quad\quad\quad\quad\quad\quad-\frac{2\mu\sqrt{2\alpha}}{1+\sqrt{\mu\alpha}}\alpha(1-L\alpha)\langle u_{t,2},\varepsilon_{t}\rangle\frac{1+\sqrt{\mu\alpha}}{2\sqrt{2\alpha}}\|z_{t}-w_{t}\|+\frac{L\alpha^{2}}{2}\|\varepsilon_{t}\|^{2}\right],$
@@
$\left.\quad\quad\quad\quad\quad\quad\quad-\frac{2\mu\sqrt{2\alpha}}{1+\sqrt{\mu\alpha}}\alpha(1-L\alpha)\langle u_{t,2},\varepsilon_{t}\rangle\frac{1+\sqrt{\mu\alpha}}{2\sqrt{2\alpha}}\|z_{t}-w_{t}\|+\frac{L\alpha^{2}}{2}\|\varepsilon_{t}\|^{2}\right]$
@@
$\left.\quad-\frac{2\mu\sqrt{2\alpha}}{1+\sqrt{\mu\alpha}}\alpha(1-L\alpha)\langle u_{t,2},\varepsilon_{t}\rangle\frac{1+\sqrt{\mu\alpha}}{2\sqrt{2\alpha}}\|z_{t}-w_{t}\|+\frac{L\alpha^{2}}{2}\|\varepsilon_{t}\|^{2}\right]+\mu\left(1+\frac{\sqrt{\mu\alpha}}{8}+\frac{2}{\sqrt{\mu\alpha}}\right)\Delta_{t}^{2}.$
@@
$\left.\left.\quad-\frac{2\mu\sqrt{2\alpha}}{1+\sqrt{\mu\alpha}}\alpha(1-L\alpha)\langle u_{t,2},\varepsilon_{t}\rangle\frac{1+\sqrt{\mu\alpha}}{2\sqrt{2\alpha}}\|z_{t}-w_{t}\|+\frac{L\alpha^{2}}{2}\|\varepsilon_{t}\|^{2}\right]+\mu\left(1+\frac{\sqrt{\mu\alpha}}{8}+\frac{2}{\sqrt{\mu\alpha}}\right)\Delta_{t}^{2}\right\}$
@@
$\left.\quad-\frac{2\mu\sqrt{2\alpha}}{1+\sqrt{\mu\alpha}}\alpha(1-L\alpha)\langle u_{t,2},\varepsilon_{t}\rangle\frac{1+\sqrt{\mu\alpha}}{2\sqrt{2\alpha}}\|z_{t}-w_{t}\|+\frac{L\alpha^{2}}{2}\|\varepsilon_{t}\|^{2}\right]$
@@
$\left.\quad\quad-\frac{2\mu\sqrt{2\alpha}}{1+\sqrt{\mu\alpha}}\alpha(1-L\alpha)\langle u_{t,2},\varepsilon_{t}\rangle\frac{1+\sqrt{\mu\alpha}}{2\sqrt{2\alpha}}\|z_{t}-w_{t}\|+\frac{L\alpha^{2}}{2}\|\varepsilon_{t}\|^{2}\right]+\frac{20\mu\Delta_{t}^{2}}{\sqrt{\mu\alpha}},$
@@
Under the same settings as in Lemma C.3 with $\phi_{t}(w)\equiv\phi(w)$, $w\in\mathbb{R}^{d}$, and $w_{t}^{*}\equiv w^{*}$, where $\phi(w)$ can be any general strongly functions in $\mathbb{R}^{d}$. We redefine $u_{t,1},u_{t,2}$ as
@@
$\quad\quad-\frac{2L\sqrt{2\alpha}}{1+\sqrt{\mu\alpha}}\alpha(1-L\alpha)\langle u_{t,2},\varepsilon_{t}\rangle\frac{1+\sqrt{\mu\alpha}}{2L\sqrt{2\alpha}}\|\nabla\phi_{t}(z_{t})-\nabla\phi_{t}(w_{t})\|+\frac{L\alpha^{2}}{2}\|\varepsilon_{t}\|^{2}.$
@@
$=\rho^{2}V_{t}-\alpha(1-L\alpha)\langle\nabla\phi_{t}(w_{t})-\nabla\phi_{t}(w^{*}),\varepsilon_{t}\rangle-\alpha(1-L\alpha)\langle\nabla\phi_{t}(z_{t})-\nabla\phi_{t}(w_{t}),\varepsilon_{t}\rangle+\frac{L\alpha^{2}}{2}\|\varepsilon_{t}\|^{2}$
@@
$\quad\quad-\frac{2L\sqrt{2\alpha}}{1+\sqrt{\mu\alpha}}\alpha(1-L\alpha)\langle u_{t,2},\varepsilon_{t}\rangle\frac{1+\sqrt{\mu\alpha}}{2L\sqrt{2\alpha}}\|\nabla\phi_{t}(z_{t})-\nabla\phi_{t}(w_{t})\|+\frac{L\alpha^{2}}{2}\|\varepsilon_{t}\|^{2}.$
@@
Under the same setting as in Lemma C.3 with $\alpha\leq\min\{1/2L,\mu/2L^{2},1/10^{3}\mu\}$, for any given $\delta\in(0,1)$ and all $t\in[T]$, the following holds with probability at least $1-\delta$ over the randomness in ${\mathcal{H}}_{t}$:
$V_{t}\leq\left(1-\frac{\sqrt{\mu\alpha}}{4}\right)^{t}V_{0}+\left(2\alpha\sigma^{2}+\frac{80\Delta^{2}}{\alpha}\right)\ln\frac{eT}{\delta}.$ (28)
@@
By Lemma C.3 we have

### v5

@@ 6 Conclusion
In this paper, we propose a new algorithm named AccBO for solving bilevel optimization problems where the upper-level is nonconvex and unbounded smooth and the lower-level problem is strongly convex. The algorithm achieved $\widetilde{O}(\epsilon^{-3})$ oracle complexity for finding an $\epsilon$-stationary point when the lower-level stochastic gradient variance is $O(\epsilon)$, which matches the rate of the state-of-the-art single-level relaxed smooth optimization [56] and is nearly optimal in terms of dependency on $\epsilon$ [2].
Limitations. There are two limitations of our work. One limitation of our work is that the convergence analysis for the Option I of our algorithm relies on the lower-level problem being a quadratic function: only under this case the algorithm becomes a single-loop procedure. Another limitation is that we require the lower-level stochastic gradient has variance $O(\epsilon)$. It remains unclear how to design single-loop algorithms for more general lower-level strongly convex functions and get rid of the small stochastic gradient variance assumption for the lower-level variable.
@@
We would like to thank the anonymous reviewers for their helpful comments. We would like to thank Tianbao Yang and Qihang Lin for helpful discussions for the quadratic function with distributional drift in the earlier version of our paper. This work has been supported by the Presidential Scholarship, the ORIEI seed funding, and the IDIA P3 fellowship from George Mason University, the Cisco Faculty Research Award, and NSF award #2436217, #2425687. The Computations were run on ARGO, a research computing cluster provided by the Office of Research Computing at George Mason University (URL: https://orc.gmu.edu).
@@ Appendix A Technical Lemmas
In this section, we will introduce a few useful lemmas. The following technical lemma on recursive control is crucial for providing high probability guarantee of the lower-level variables $y_{t}$ and $\hat{y}_{t}$ in Algorithm 2 at anytime. We follow a similar argument as in [20, Proposition 29] with a slight generalization.
@@ Appendix B Auxiliary Lemmas for Bilevel Optimization
Lemma B.1 (Lipschitz property, [38, Lemma 8]).
@@
Lemma B.2 ($(L_{0},L_{1})$-smoothness, [38, Lemma 9]).
@@
Lemma B.3 (Descent inequality, [38, Lemma 10]).
@@
$\left.-\frac{2\mu\sqrt{2\alpha}}{1+\sqrt{\mu\alpha}}\alpha(1-L\alpha)\langle u_{t,2},\varepsilon_{t}\rangle\frac{1+\sqrt{\mu\alpha}}{2\sqrt{2\alpha}}\|z_{t}-w_{t}\|+\frac{\alpha(1+L\alpha)}{2}\|\varepsilon_{t}\|^{2}\right]+\frac{20\mu\Delta_{t}^{2}}{\sqrt{\mu\alpha}}.$
@@
Let us define vector $\omega_{t}\in\mathbb{R}^{d}$ and matrices ${\mathbf{A}}\in\mathbb{R}^{2d\times 2d},{\mathbf{B}}\in\mathbb{R}^{2d\times d}$ as
$\omega_{t}=\nabla\phi_{t}(z_{t}),\quad{\mathbf{A}}=\begin{bmatrix}1+\gamma&-\gamma\\ 1&0\end{bmatrix}\otimes{\mathbf{I}}_{d},\quad{\mathbf{B}}=\begin{bmatrix}-\alpha\\ 0\end{bmatrix}\otimes{\mathbf{I}}_{d}.$
By (9) we have
${}^{\top}={\mathbf{A}}\theta_{t}+{\mathbf{B}}(\omega_{t}+\varepsilon_{t})$ (14)
Since $\phi_{t}$ is $\mu$-strongly convex, then
@@
where matrix ${\mathbf{X}}_{1}\in\mathbb{R}^{3d\times 3d}$ is defined as
${\mathbf{X}}_{1}=\frac{1}{2}\begin{bmatrix}\mu\gamma^{2}&-\mu\gamma^{2}&-\gamma\\ -\mu\gamma^{2}&\mu\gamma^{2}&\gamma\\ -\gamma&\gamma&\alpha(2-L\alpha)\\ \end{bmatrix}\otimes{\mathbf{I}}_{d}.$
@@
where matrix ${\mathbf{X}}_{2}\in\mathbb{R}^{3d\times 3d}$ is defined as
${\mathbf{X}}_{2}=\frac{1}{2}\begin{bmatrix}\mu(1+\gamma)^{2}&-\mu\gamma(1+\gamma)&-(1+\gamma)\\ -\mu\gamma(1+\gamma)&\mu\gamma^{2}&\gamma\\ -(1+\gamma)&\gamma&\alpha(2-L\alpha)\\ \end{bmatrix}\otimes{\mathbf{I}}_{d}.$
@@
By [42, Section 3.1] (see also [5, Corollary 4.9], [4, Theorem 2.3]), we have the following fact:
$\begin{bmatrix}{\mathbf{A}}^{\top}{\mathbf{P}}{\mathbf{A}}-\rho^{2}{\mathbf{P}}&{\mathbf{A}}^{\top}{\mathbf{P}}{\mathbf{B}}\\ {\mathbf{B}}^{\top}{\mathbf{P}}{\mathbf{A}}&{\mathbf{B}}^{\top}{\mathbf{P}}{\mathbf{B}}\end{bmatrix}-(\rho^{2}{\mathbf{X}}_{1}+(1-\rho^{2}){\mathbf{X}}_{2})\preceq 0,$ (21)
which, combined with (20) yields
$\begin{bmatrix}w_{t+1}-w_{t}^{*}\\ w_{t}-w_{t}^{*}\end{bmatrix}^{\top}{\mathbf{P}}\begin{bmatrix}w_{t+1}-w_{t}^{*}\\ w_{t}-w_{t}^{*}\end{bmatrix}-\rho^{2}\theta_{t}^{\top}{\mathbf{P}}\theta_{t}$
$\quad=\begin{bmatrix}\theta_{t}\\ \omega_{t}\end{bmatrix}^{\top}\begin{bmatrix}{\mathbf{A}}^{\top}{\mathbf{P}}{\mathbf{A}}-\rho^{2}{\mathbf{P}}&{\mathbf{A}}^{\top}{\mathbf{P}}{\mathbf{B}}\\ {\mathbf{B}}^{\top}{\mathbf{P}}{\mathbf{A}}&{\mathbf{B}}^{\top}{\mathbf{P}}{\mathbf{B}}\end{bmatrix}\begin{bmatrix}\theta_{t}\\ \omega_{t}\end{bmatrix}+\varepsilon_{t}^{\top}{\mathbf{B}}^{\top}{\mathbf{P}}{\mathbf{B}}\varepsilon_{t}$
$\quad\leq\begin{bmatrix}\theta_{t}\\ \omega_{t}\end{bmatrix}^{\top}(\rho^{2}{\mathbf{X}}_{1}+(1-\rho^{2}){\mathbf{X}}_{2})\begin{bmatrix}\theta_{t}\\ \omega_{t}\end{bmatrix}+\frac{\alpha}{2}\|\varepsilon_{t}\|^{2}$
$\quad\leq-(\phi_{t}(w_{t+1})-\phi_{t}(w_{t}^{*}))+\rho^{2}(\phi_{t}(w_{t})-\phi_{t}(w_{t}^{*}))-\alpha(1-L\alpha)\langle\nabla\phi_{t}(z_{t}),\varepsilon_{t}\rangle+\frac{\alpha+L\alpha^{2}}{2}\|\varepsilon_{t}\|^{2},$
where the first inequality uses (14) and (21), along with the fact that ${\mathbf{B}}^{\top}{\mathbf{P}}{\mathbf{B}}=(\alpha/2)\otimes{\mathbf{I}}_{d}$ and hence $\lambda_{\max}({\mathbf{B}}^{\top}{\mathbf{P}}{\mathbf{B}})=\alpha/2$; and the last inequality follows by (20). Rearrange the above inequality and by definition of the potential function $V_{t}$, we obtain
@@
$\leq\rho^{2}V_{t}-\alpha(1-L\alpha)\langle\nabla\phi_{t}(z_{t}),\varepsilon_{t}\rangle+\frac{\alpha(1+L\alpha)}{2}\|\varepsilon_{t}\|^{2}.$
@@
$\quad\quad\leq\rho^{2}V_{t}-\mu\alpha(1-L\alpha)\langle z_{t}-w_{t}^{*},\varepsilon_{t}\rangle+\frac{\alpha+L\alpha^{2}}{2}\|\varepsilon_{t}\|^{2}$
$\quad\quad=\rho^{2}V_{t}-\mu\alpha(1-L\alpha)\langle w_{t}-w_{t}^{*},\varepsilon_{t}\rangle-\mu\alpha(1-L\alpha)\langle z_{t}-w_{t},\varepsilon_{t}\rangle+\frac{\alpha(1+L\alpha)}{2}\|\varepsilon_{t}\|^{2}$
@@
$\quad\quad\quad\quad\quad\quad-\frac{2\mu\sqrt{2\alpha}}{1+\sqrt{\mu\alpha}}\alpha(1-L\alpha)\langle u_{t,2},\varepsilon_{t}\rangle\frac{1+\sqrt{\mu\alpha}}{2\sqrt{2\alpha}}\|z_{t}-w_{t}\|+\frac{\alpha(1+L\alpha)}{2}\|\varepsilon_{t}\|^{2},$
@@
$\left.\quad\quad\quad\quad\quad-\frac{2\mu\sqrt{2\alpha}}{1+\sqrt{\mu\alpha}}\alpha(1-L\alpha)\langle u_{t,2},\varepsilon_{t}\rangle\frac{1+\sqrt{\mu\alpha}}{2\sqrt{2\alpha}}\|z_{t}-w_{t}\|+\frac{\alpha(1+L\alpha)}{2}\|\varepsilon_{t}\|^{2}\right]$
@@
$\left.\quad\quad\quad\quad\quad-\frac{2\mu\sqrt{2\alpha}}{1+\sqrt{\mu\alpha}}\alpha(1-L\alpha)\langle u_{t,2},\varepsilon_{t}\rangle\frac{1+\sqrt{\mu\alpha}}{2\sqrt{2\alpha}}\|z_{t}-w_{t}\|+\frac{\alpha(1+L\alpha)}{2}\|\varepsilon_{t}\|^{2}\right],$
@@
$\left.\quad\quad\quad\quad\quad-\frac{2\mu\sqrt{2\alpha}}{1+\sqrt{\mu\alpha}}\alpha(1-L\alpha)\langle u_{t,2},\varepsilon_{t}\rangle\frac{1+\sqrt{\mu\alpha}}{2\sqrt{2\alpha}}\|z_{t}-w_{t}\|+\frac{\alpha(1+L\alpha)}{2}\|\varepsilon_{t}\|^{2}\right]$
@@
$\left.\quad-\frac{2\mu\sqrt{2\alpha}}{1+\sqrt{\mu\alpha}}\alpha(1-L\alpha)\langle u_{t,2},\varepsilon_{t}\rangle\frac{1+\sqrt{\mu\alpha}}{2\sqrt{2\alpha}}\|z_{t}-w_{t}\|+\frac{\alpha(1+L\alpha)}{2}\|\varepsilon_{t}\|^{2}\right]+\mu\left(1+\frac{\sqrt{\mu\alpha}}{8}+\frac{2}{\sqrt{\mu\alpha}}\right)\Delta_{t}^{2}.$
@@
$\left.\left.\quad-\frac{2\mu\sqrt{2\alpha}}{1+\sqrt{\mu\alpha}}\alpha(1-L\alpha)\langle u_{t,2},\varepsilon_{t}\rangle\frac{1+\sqrt{\mu\alpha}}{2\sqrt{2\alpha}}\|z_{t}-w_{t}\|+\frac{\alpha(1+L\alpha)}{2}\|\varepsilon_{t}\|^{2}\right]+\mu\left(1+\frac{\sqrt{\mu\alpha}}{8}+\frac{2}{\sqrt{\mu\alpha}}\right)\Delta_{t}^{2}\right\}$
@@
$\left.\quad-\frac{2\mu\sqrt{2\alpha}}{1+\sqrt{\mu\alpha}}\alpha(1-L\alpha)\langle u_{t,2},\varepsilon_{t}\rangle\frac{1+\sqrt{\mu\alpha}}{2\sqrt{2\alpha}}\|z_{t}-w_{t}\|+\frac{\alpha(1+L\alpha)}{2}\|\varepsilon_{t}\|^{2}\right]$
@@
$\left.\quad-\frac{2\mu\sqrt{2\alpha}}{1+\sqrt{\mu\alpha}}\alpha(1-L\alpha)\langle u_{t,2},\varepsilon_{t}\rangle\frac{1+\sqrt{\mu\alpha}}{2\sqrt{2\alpha}}\|z_{t}-w_{t}\|+\frac{\alpha(1+L\alpha)}{2}\|\varepsilon_{t}\|^{2}\right]+\frac{20\mu\Delta_{t}^{2}}{\sqrt{\mu\alpha}},$
@@
Under the same settings as in Lemma C.3 with $\phi_{t}(w)\equiv\phi(w)$, and $w_{t}^{*}\equiv w^{*}$, where $\phi(w)$ can be any general strongly functions in $\mathbb{R}^{d}$. We redefine $u_{t,1},u_{t,2}$ as
@@
$-\frac{2L\sqrt{2\alpha}}{1+\sqrt{\mu\alpha}}\alpha(1-L\alpha)\langle u_{t,2},\varepsilon_{t}\rangle\frac{1+\sqrt{\mu\alpha}}{2L\sqrt{2\alpha}}\|\nabla\phi_{t}(z_{t})-\nabla\phi_{t}(w_{t})\|+\frac{\alpha(1+L\alpha)}{2}\|\varepsilon_{t}\|^{2}.$
@@
$=\rho^{2}V_{t}-\alpha(1-L\alpha)\langle\nabla\phi_{t}(w_{t})-\nabla\phi_{t}(w^{*}),\varepsilon_{t}\rangle-\alpha(1-L\alpha)\langle\nabla\phi_{t}(z_{t})-\nabla\phi_{t}(w_{t}),\varepsilon_{t}\rangle+\frac{\alpha(1+L\alpha)}{2}\|\varepsilon_{t}\|^{2}$
@@
$\quad\quad-\frac{2L\sqrt{2\alpha}}{1+\sqrt{\mu\alpha}}\alpha(1-L\alpha)\langle u_{t,2},\varepsilon_{t}\rangle\frac{1+\sqrt{\mu\alpha}}{2L\sqrt{2\alpha}}\|\nabla\phi_{t}(z_{t})-\nabla\phi_{t}(w_{t})\|+\frac{\alpha(1+L\alpha)}{2}\|\varepsilon_{t}\|^{2}.$
@@
Under the same setting as in Lemma C.3 with $\alpha\leq 1/25L$, for any given $\delta\in(0,1)$ and all $t\in[T]$, the following holds with probability at least $1-\delta$ over the randomness in ${\mathcal{H}}_{t}$:
$V_{t}\leq\left(1-\frac{\sqrt{\mu\alpha}}{4}\right)^{t}V_{0}+\left(\frac{5\sqrt{\alpha}\sigma^{2}}{\sqrt{\mu}}+\frac{80\Delta^{2}}{\alpha}\right)\ln\frac{eT}{\delta}.$ (30)
@@
By Lemma C.3 and the choice of $\alpha\leq 1/L$, we have